OpenStudy (anonymous):

Jessica deposits $300 into a savings account that pays an annual interest rate of 2%, compounded twice a year. How much money will Jessica have in her account at the end of one year?

OpenStudy (amistre64):

it pays 2% interest per year, but it does it in 2 installments; 1% and 1%, it compounds the first result into the second so you receive a little more in the end.

OpenStudy (anonymous):

i always forget this formula is it 300( 1 + .01)^2 ?

OpenStudy (amistre64):

(300(1.01))(1.01) is what im seeing,

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

A = P(1+r/2)^2

OpenStudy (amistre64):

well, generically, which is what i meant to type: A = P(1+r/n)^nt where P = starting amount, r = rate, n=compounded cycle, and t = years

OpenStudy (anonymous):

yes - thats it - i have trouble remembering that one - id better make it into a mantra!!

OpenStudy (anonymous):

so amount comes to 306.3

OpenStudy (anonymous):

sorry 306.03

OpenStudy (anonymous):

Thank You guys very much appreciated